# FRAP: Faithful and Realistic Text-to-Image Generation with Adaptive Prompt Weighting

### v1

@@ Introduction
Prior works (Chefer et al. 2023; Li et al. 2023) improve the semantic guidance in SD by optimizing the latent code during inference. These methods demonstrate excellent capability in addressing semantic compliance issues. However, the latent code could go out-of-distribution (OOD) after several iterative updates, which could lead to generation of unrealistic images.
To address this issue, we propose a simple, yet effective approach based on the prompt weighting technique that improves faithfulness while maintaining the realistic appearance of generated images. Unlike the manual approach, FRAP adaptively updates the weight coefficients of different text tokens during the RGP to emphasize or de-emphasize certain tokens in order to facilitate improving faithfulness. The weights are guided by our unified objective function defined on the CA maps, strengthening the object presence and encouraging object-modifier binding.
The following subsections describe our proposed unified objective function and adaptive prompt weighting method. The overview of FRAP is illustrated in Fig. 1. In the supplementary materials, we present the overall algorithm and implementation details of our proposed method.
Unified Objective
We propose a novel objective function $\mathcal{L}$, defined on the CA maps, to address improving the alignment of a generated image with the prompt. This unified objective function is used as the loss for on-the-fly optimization of the adaptive prompt weight parameters. Our proposed unified objective function $\mathcal{L}_{t}$ at time-step $t$ consists of two terms: (i) the object presence loss $\mathcal{L}_{\text{presence,t}}$ and (ii) the object-modifier binding loss $\mathcal{L}_{\text{binding},t}$. The overall objective function is defined as:
@@
The object presence loss is defined on the set of object tokens $S$ in the text prompt to reinforce the presence of each object $s\in S$ in the generated image to prevent neglect or generating hybrid creatures. The object-modifier binding loss is defined on the set of object-modifier token pairs $(s,r)\in P(s,R_{s})$ for $s\in S$, to encourage the correct binding of each object token $s$ and each of its related modifier token $r\in R_{s}$. To automate the manual identification of object and modifier tokens in (Chefer et al. 2023; Li et al. 2023), we follow SynGen (Rassin et al. 2023) and use the spaCy (Honnibal and Montani 2017) language parser to extract the object and modifier tokens as explained in the supplementary materials.
Object Presence Loss.
@@
Object-Modifier Binding Loss.
@@
$\mathcal{L}_{\text{binding},t}^{s,r}=\frac{1}{P^{2}}\sum_{\mathrm{pixels}}\Bigg{(}\mathrm{min}\Big{(}\mathrm{Pr}\big{(}G(A_{t}^{s})\big{)},\mathrm{Pr}\big{(}G_{a}(A_{t}^{r})\big{)}\Big{)}\Bigg{)},$ (6)
where $P$ is the spatial dimension of the CA map, $G_{a}(A_{t}^{r})$ denotes the aligned smoothed CA map of the token $r$ according to token $s$ following (Li et al. 2023), and $\mathrm{Pr}\big{(}G(A_{t}^{s})\big{)}$ is the normalized discrete probability distribution of the smoothed CA map of token $s$. This distribution is obtained by applying Softmax over all pixels of the respective CA map. We apply the same process to get $\mathrm{Pr}\big{(}G_{a}(A_{t}^{r})\big{)}$.
Next, as shown in Eq. (4), we maximize this minimum overlap between the two discrete probability distributions in a minimax fashion.555 Note the negative sign for $\mathcal{L}_{\text{binding},t}$ in Eq. (4), which results in maximizing the minimum overlap. The minmax approach aims to ensure a robust match between the CA maps. This is particularly useful for representing small objects in a scene. Moreover, it is less susceptible to the influence of outliers compared to divergence methods such as JSD, which is beneficial in handling noisy and uncertain CA maps.
Overall Loss.
We calculate the total presence loss as the mean of per-token object presence loss for all object tokens $s\in S$:
@@
Note that our per-token presence loss in Eq. (5) is the same as A&E’s (Chefer et al. 2023). However, different from A&E, we use the mean among all objects as the total object presence loss instead of only using the loss from the most neglected token with the lowest max activation. This enhances the presence of all objects instead of a single one, which is expected to be essential for complex prompts. We ablate our loss design in the supplementary materials.
Adaptive Prompt Weighting
Different from the manual prompt weighting (Face 2023a; Stewart 2023), we propose FRAP, an adaptive prompt weighting method, to update the weight coefficients during the RGP automatically. We update the per-token weight coefficients $\phi$ after each time-step $t$ according to our unified loss objective $\mathcal{L}_{t}$ (see Eq. (4)) defined on the internal feedback from the CA maps. Our approach determines the per-token weights automatically and appropriately for each token to encourage the SD (Rombach et al. 2022) model to generate images that are faithful to the semantics and have a realistic appearance.
@@
At time-step $t\!=\!T$, we initialize $\alpha_{T}\!=\!\vec{0}$ so we start from a trivial weighting of $\phi_{T}\!=\!\vec{1}$ (equivalent to the standard SD with no prompt weighting). Then, our on-the-fly optimization will gradually adjust the weight coefficients to improve alignment with the prompt. This optimization step takes place during inference with no additional overhead for model training. We only perform one UNet inference call at each time-step, and we use the weight parameter $\alpha_{t}$ obtained from the optimization in the previous time-step $t+1$ to get the current weight coefficient $\phi_{t}$. Thus, our adaptive prompt weighting process keeps the same number of UNet calls as the standard SD (Rombach et al. 2022). In contrast, other methods such as A&E (Chefer et al. 2023) and D&B (Li et al. 2023) perform several optimization refinement steps at each time-step.
Experiments
@@
Table 1: Quantitative comparison of FRAP and other methods on three complex datasets Color-Object-Scene, COCO-Subject, and COCO-Attribute. Higher is better (except for latency). Our method is efficient, for instance, FRAP is on average $4$ and $5$ seconds faster than D&B and A&E respectively, for generating one image on the COCO-Subject dataset.
@@
Prompt-Image Alignment
Image Quality Assesment
@@
Color-Obj-Scene  (Li et al. 2023) SD (Rombach et al. 2022) 0.707 0.366 0.254 0.282 0.675 0.770 8.0
A&E (Chefer et al. 2023) 0.708 0.376 0.267 0.286 0.675 0.770 17.9
D&B (Li et al. 2023) 0.719 0.375 0.264 0.286 0.682 0.808 16.7
FRAP (ours) 0.727 0.381 0.269 0.287 0.676 0.835 14.3
COCO-Subject  (Li et al. 2023) SD (Rombach et al. 2022) 0.829 0.331 0.250 0.278 0.805 0.852 7.9
A&E (Chefer et al. 2023) 0.829 0.333 0.254 0.277 0.805 0.852 19.1
D&B (Li et al. 2023) 0.835 0.333 0.253 0.279 0.814 0.862 18.1
FRAP (ours) 0.837 0.334 0.255 0.279 0.814 0.866 14.0
COCO-Attribute  (Li et al. 2023) SD (Rombach et al. 2022) 0.817 0.347 0.245 0.286 0.703 0.750 8.1
A&E (Chefer et al. 2023) 0.820 0.353 0.254 0.286 0.703 0.750 15.3
D&B (Li et al. 2023) 0.821 0.353 0.251 0.288 0.720 0.760 16.7
FRAP (ours) 0.838 0.358 0.254 0.289 0.725 0.778 13.9

### v2

@@ 4 Method
Prior works (Chefer et al., 2023; Li et al., 2023) improve the semantic guidance in SD by optimizing the latent code during inference. These methods demonstrate excellent capability in addressing semantic compliance issues. However, the latent code could go out-of-distribution (OOD) after several iterative updates, which could lead to generation of unrealistic images.
To address this issue, we propose a simple, yet effective approach based on the prompt weighting technique that improves faithfulness while maintaining the realistic appearance of generated images. Unlike the manual approach, FRAP adaptively updates the weight coefficients of different text tokens during the RGP to emphasize or de-emphasize certain tokens in order to facilitate improving faithfulness. The weight updates are guided by our unified objective function defined on the CA maps, strengthening the object presence and encouraging object-modifier binding.
The following subsections describe our proposed unified objective function and adaptive prompt weighting method. The overview of FRAP is illustrated in Fig. 1. In App. A, we present the overall algorithm and implementation details of our proposed method. We include our code in the supplementary material and will open-source our codebase for further research after review.
4.1 Unified Objective
We propose an objective function $\mathcal{L}$, formulated based on the CA maps, to enhance the alignment between the generated image and the prompt. This unified objective function is used as the loss for on-the-fly optimization of the adaptive prompt weight parameters. Our proposed unified objective function $\mathcal{L}_{t}$ at time-step $t$ consists of two terms: (i) the object presence loss $\mathcal{L}_{\text{presence,t}}$ and (ii) the object-modifier binding loss $\mathcal{L}_{\text{binding},t}$. The overall objective function is defined as:
@@ 4.1 Unified Objective
To automate the manual identification of object tokens and modifier tokens from the prompt, we follow SynGen (Rassin et al., 2023) and use the spaCy (Honnibal & Montani, 2017) language parser to extract the set $S$ of all object tokens $s\in S$ and set $R$ of all modifier tokens $r\in R$ in the prompt. For each object token $s$, we use $R_{s}$ to represent its related modifier tokens $r\in R_{s}$. We use $(s,r)$ to represent an object-modifier pair (e.g. “brown dog”), where $P$ is the pairing operation and $P(s,R_{s})$ represents the set of all object-modifier pairs between object token $s$ and its related modifier tokens $r\in R_{s}$. We use $P(S,R)$ to represents the set of all object-modifier pairs identified in the prompt. The object presence loss is defined on the set of object tokens $S$ in the text prompt to reinforce the presence of each object $s\in S$ in the generated image to prevent neglect or generating hybrid creatures. The object-modifier binding loss is defined on the set of all object-modifier token pairs $(s,r)\in P(S,R)$ to encourage the correct binding of each object token $s\in S$ and each of its related modifier token $r\in R_{s}$. More details on object and modifier token identification are explained in App. A.3.
4.1.1 Object Presence Loss
@@
As pointed out in the ablation studies done by Chefer et al. (2023), the Gaussian smoothing applied to the CA map is important, since it ensures that maximizing the largest activation value not only encourages a single point of high activation but rather a region of high activation values, which is critical for generating the entire object and avoiding partial generation of objects.
Then, we calculate the total presence loss as the mean of per-token object presence loss for all object tokens $s\in S$:
Note that the per-token presence loss in Eq. (5) is adopted from A&E’s (Chefer et al., 2023) design. However, our design of the total object presence loss in Eq. (6) is original and different from A&E. We use the mean among all objects as the total object presence loss instead of only using the loss from the most neglected token with the lowest max activation. Our design enhances the presence of all objects instead of a single one, which is expected to be essential for complex prompts. Through our empirical evaluation in Table 6 in App. C, our design which enhances the presence of all objects outperforms A&E’s design where only the most neglected token is considered.
We also consider an alternative design for the object presence loss, where we attempt to replace the object presence loss with the Total Variation (TV) loss adopted by D&B (Li et al., 2023), which encourages high activation difference across spatial locations. Through our empirical evaluation in Table 7 in App. C, the object presence loss in Eq. (5) performs better than the TV loss.
4.1.2 Object-Modifier Binding Loss
@@
$\mathcal{L}_{\text{binding},t}^{s,r}=\frac{1}{L^{2}}\sum_{\mathrm{pixels}}\Bigg{(}\mathrm{min}\Big{(}\mathrm{Pr}\big{(}G(A_{t}^{s})\big{)},\mathrm{Pr}\big{(}G_{a}(A_{t}^{r})\big{)}\Big{)}\Bigg{)},$ (7)
where $L$ is the spatial dimension of the CA map, $G_{a}(A_{t}^{r})$ denotes the aligned smoothed CA map of the token $r$ according to token $s$ following Li et al. (2023), and $\mathrm{Pr}\big{(}G(A_{t}^{s})\big{)}$ is the normalized discrete probability distribution of the smoothed CA map of token $s$. This distribution is obtained by applying Softmax over all pixels of the respective CA map. We apply the same process to get $\mathrm{Pr}\big{(}G_{a}(A_{t}^{r})\big{)}$.
@@
As shown in Eq. (4), we maximize this minimum overlap between the two discrete probability distributions in a minimax fashion.555 Note the negative sign for $\mathcal{L}_{\text{binding},t}$ in Eq. (4), which results in maximizing the minimum overlap. Our minimax approach aims to ensure that the least probable events have high probabilities in both distributions. This guarantees that every pixel of their respective object or modifier receives a considerable probability, regardless of which CA map is more accurate overall. This is particularly useful in cases where there is a need to depict a small object in the scene. We found this approach to be more beneficial than Minimizing JSD (as in D&B (Li et al., 2023)) or KLD (as in SynGen (Rassin et al., 2023)), likely due to the following two reasons: (i) it ensures a robust match between the CA maps. This is particularly useful for representing small objects in a scene. (ii) it is less susceptible to the influence of outliers compared to divergence methods such as JSD and KLD. This robustness is beneficial in handling noisy and uncertain CA maps. Through our empirical evaluation in Table 8 in App. C, our object-modifier binding loss in Eq. (7) performs better than either the JSD loss or the KLD loss.
4.2 Adaptive Prompt Weighting
Different from the manual prompt weighting (Hugging Face, 2023a; Stewart, 2023), we propose FRAP, an adaptive prompt weighting method, to update the weight coefficients during the RGP automatically. We update the per-token weight coefficients $\phi$ after each time-step $t$ according to our unified loss objective $\mathcal{L}_{t}$ (see Eq. (4)) defined on the internal feedback from the CA maps. Our approach determines the per-token weights automatically and appropriately for each token to encourage the SD (Rombach et al., 2022) model to generate images that are faithful to the semantics and have a realistic appearance.
@@ 4.2 Adaptive Prompt Weighting
At time-step $t\!=\!T$, we initialize $\alpha_{T}\!=\!\vec{0}$ so we start from a trivial weighting of $\phi_{T}\!=\!\vec{1}$ (equivalent to the vanilla SD with no prompt weighting). Then, our on-the-fly optimization will gradually adjust the weight coefficients to improve alignment with the prompt. This optimization step takes place during inference with no additional overhead for model training. We only perform one UNet inference call at each time-step, and we use the weight parameter $\alpha_{t}$ obtained from the optimization in the previous time-step $t+1$ to get the current weight coefficient $\phi_{t}$. Thus, our adaptive prompt weighting process keeps the same number of UNet calls as the vanilla SD (Rombach et al., 2022). In contrast, other methods such as A&E (Chefer et al., 2023) and D&B (Li et al., 2023) perform several optimization refinement steps at each time-step.
5 Experiments
In this section, we provide quantitative evaluation and visual comparisons of the effectiveness of FRAP in terms of prompt-image alignment and image quality assessment. We also assess the efficiency of FRAP. In addition, we apply FRAP to the recent prompt rewrite method. Moreover, we perform human evaluations in addition to using automated evaluation metrics.
In App. C, we present an ablation study on the design choices of FRAP, including a comparison to alternative loss functions. In App. D, we perform evaluations on additional model, baselines, and even larger datasets. Moreover, we provide additional visual comparisons in Figs. 13 and 14 in the appendix.
Experimental Settings. We compare our results to SD (Rombach et al., 2022) and to recent methods A&E (Chefer et al., 2023) and D&B (Li et al., 2023). All baselines and our approach are evaluated based on the SDv1.5 (Rombach et al., 2022; Hugging Face, 2023b) model to align with D&B (Li et al., 2023).666 Note, however, that FRAP is also applicable to SDXL (Podell et al., 2023) and improves its performance, as evident by our additional experiments (see Table 15 in App. D). We evaluate on three Simple, manually crafted prompt datasets from A&E: Animal-Animal (S-AA), Color-Object (S-CO), and Animal-Object (S-AO); and five Complex datasets from D&B: Animal-Scene (C-AS), Color-Object-Scene (C-COS), Multi-Object (C-MO), COCO-Attribute (C-CA), and COCO-Subject (C-CS). For comparing the efficiency of different methods, we measure the end-to-end latency and total number of UNet calls. In the experiments, our method utilizes the unified objective function on early time-step cross-attention maps to efficiently select an initial latent code with decent object presence and object-modifier binding. We obtain the initial latent code by performing 15 initial inference steps on a batch of 4 noisy latent code samples and select one with minimal loss. In App. A, we further elaborate on the baselines, implementation details, prompt datasets, evaluation metrics, and other experimental settings.
5.1 Prompt-Image Alignment, Overall Image Quality, and Image Authenticity
We use multiple metrics to evaluate the performance of various methods in three different aspects: prompt-image alignment, overall image quality, and image authenticity. For prompt-image alignment, we evaluate metrics including Text-to-Text Similarity (TTS) (Chefer et al., 2023; Li et al., 2023), Full-prompt Similarity (Full-CLIP) (Chefer et al., 2023), and Minimum Object Similarity (MOS) (Chefer et al., 2023). Each metric reflects how well the generated image aligns with the given prompt at a different semantic level. TTS (Chefer et al., 2023; Li et al., 2023) evaluates the sentence-to-sentence level alignment, using the BLIP (Li et al., 2022) captioning model to generate a caption for the synthesized image and compares the cosine similarity between the CLIP (Radford et al., 2021) embeddings of the original prompt and the generated caption of the synthesized image. Full-CLIP (Chefer et al., 2023) reflects the sentence-to-image level alignment by measuring the cosine similarity between CLIP embeddings of the original text prompt and CLIP embeddings of the synthesized image. MOS (Chefer et al., 2023) considers a finer-grained alignment between the sub-sentence (i.e. part of the prompt) and the synthesized image. To evaluate the overall image quality (i.e. image sharpness, brightness), we utilize the Human Preference Score version 2 (HPSv2) (Wu et al., 2023) which is trained on human preference datasets and the CLIP-IQA (Wang et al., 2023) metric which is overall CLIP-IQA score averaged over four criteria including “quality”, “noisiness”, “natural”, and “real”. For image authenticity, we adopt the “real” component of the CLIP-IQA metric (i.e. CLIP-IQA-Real) which specifically reflects the level of image authenticity and how realistic the generated images are. See App. A.3 for more detailed description of the evaluation metrics.
We have placed the detailed performance over all datasets and evaluation metrics in the larger Table 5 in App. B. Summarizing the results here, we observe a tight competition among A&E, D&B, and FRAP across all metrics in datasets with simple hand-crafted templates (S-AA, S-CO, and S-AO), while all three methods exceed the performance of vanilla SD. This shows that the simple datasets may not be sufficient to definitively identify the superior method among the competing options due to the simplicity of the prompts.
We see a clear advantage of FRAP on more challenging complex datasets. In Table 1, FRAP achieves either comparable or improved overall image quality in terms of HPSv2 and CLIP-IQA while having a lower latency than A&E and D&B. More importantly, FRAP achieves significant gain in CLIP-IQA-Real score on the Color-Obj-Scene, COCO-Attribute datasets and consistently achieves the highest performance in image authenticity reflected by the CLIP-IQA-Real metric. Whereas the prior methods could potentially drive the latent code out-of-distribution (OOD) causing the low image authenticity reflected by lower CLIP-IQA-Real scores despite having decent prompt-image alignment and overall image quality (see Fig. 3 where images generated by A&E and D&B have cartoonish appearances in “turtle”, “horse”, and “clock”). We also updated Table 1 to include statistical significance t-test results. For prompt-image alignment, FRAP significantly improves TTS score and Full-CLIP score on the Color-Obj-Scene and COCO-Attribute datasets which are the more challenging datasets since they include modifier words to the object words, the improvement on these two datasets signifying that FRAP handles object-modifier binding well. Whereas FRAP achieves comparable prompt-image alignment on the simpler COCO-Subject dataset which only contains object words without any modifiers, making it easier for all methods to only deal with object presence without handling any object-modifier binding. We further discuss the observed limitations in App. E, where all evaluated methods (FRAP and all other methods including Standard SD (Rombach et al., 2022), A&E (Chefer et al., 2023), and D&B (Li et al., 2023)) fail to generate the correct number of objects on the Multi-Object dataset (e.g., the count of generated objects does not match the prompt).
@@ 5.1 Prompt-Image Alignment, Overall Image Quality, and Image Authenticity
Table 1: Quantitative comparison of FRAP and other methods on three complex datasets Color-Object-Scene, COCO-Attribute, and COCO-Subject. Higher is better (except for latency). Our method is efficient, for instance, FRAP is on average $4$ and $5$ seconds faster than D&B and A&E respectively, for generating one image on the COCO-Subject dataset. * means that FRAP’s improvement over the next best method is statistically significant at $p<0.05$ with a t-test.
@@
Prompt-Image
Alignment
Overall
Image Quality
Image
Authenticity
@@
Color-Obj-Scene  (Li et al., 2023) SD (2022) 0.707 0.366 0.254 0.282 0.675 0.770 8.0
A&E (2023) 0.708 0.376 0.267 0.286 0.675 0.770 17.9
D&B (2023) 0.719 0.375 0.264 0.286 0.682 0.808 16.7
FRAP (ours) 0.727* 0.381* 0.269* 0.287 0.676 0.835* 14.3
COCO-Attribute  (Li et al., 2023) SD (2022) 0.817 0.347 0.245 0.286 0.703 0.750 8.1
A&E (2023) 0.820 0.353 0.254 0.286 0.703 0.750 15.3
D&B (2023) 0.821 0.353 0.251 0.288 0.720 0.760 16.7
FRAP (ours) 0.838* 0.358* 0.254 0.289 0.725 0.778* 13.9
COCO-Subject  (Li et al., 2023) SD (2022) 0.829 0.331 0.250 0.278 0.805 0.852 7.9
A&E (2023) 0.829 0.333 0.254 0.277 0.805 0.852 19.1
D&B (2023) 0.835 0.333 0.253 0.279 0.814 0.862 18.1
FRAP (ours) 0.837 0.334* 0.255 0.279 0.814 0.866 14.0
In addition, we evaluate the FID (Fréchet Inception Distance) metric (Heusel et al., 2017) which is a full-reference IQA metric that requires a ground-truth reference image in addition to the generated image, different from the no-reference IQA metrics (i.e. HPSv2 (Wu et al., 2023) and CLIP-IQA (Wang et al., 2023)) which only requires the generated image. Therefore, we follow the evaluation settings of prior works in text-to-image generation (Rombach et al., 2022; Saharia et al., 2022; Ramesh et al., 2022; Podell et al., 2023) and adopt the validation set of the MS-COCO dataset (Lin et al., 2014) which contains the ground-truth reference image. To select prompt-image pairs that are most relevant to prompt-image alignment evaluation and specifically object presence and object-modifier binding, we only keep the relevant prompts with at least two objects and each has at least one modifier word by filtering the prompts with the spaCy (Honnibal & Montani, 2017) language dependency parser. This filtering process selects 16k most relevant prompts from the original 40k prompts in MS-COCO, and we randomly sample a 5k subset from the 16k most relevant prompts. We refer to this dataset as COCO-5K and will release this dataset to facilitate reproducibility and further research.
In Table 2, we observe significant improvements and best performance for FID (the lower the FID, the better) and all other evaluated metrics for prompt-image alignment, overall image quality, and image authenticity for the COCO-5K dataset. Especially, there is a substantial improvement in CLIP-IQA-Real score over the three other evaluated methods showing the advantage of achieving high image authenticity even when optimizing for prompt-image alignment when using our proposed FRAP method.
Table 2: Quantitative comparison of FRAP and other methods on the COCO-5K subset from the MS-COCO (Lin et al., 2014) dataset. Higher is better (except for FID and latency). Our method is efficient, for instance, FRAP is on average $7$ and $10.1$ seconds faster than D&B and A&E respectively, for generating one image on the COCO-5K dataset. * means that FRAP’s improvement over the next best method is statistically significant at $p<0.05$ with a t-test (note that the t-test does not apply to FID since it is a distance measure between two distributions with no standard deviation).
Method FID $\downarrow$
Prompt-Image
Alignment
Overall
Image Quality
Image
Authenticity
Latency (s) $\downarrow$
TTS $\uparrow$
Full-
CLIP $\uparrow$
MOS $\uparrow$
HPS
v2 $\uparrow$
CLIP-
IQA $\uparrow$
CLIP-IQA
Real $\uparrow$
SD (2022) 39.52 0.801 0.322 0.247 0.276 0.712 0.784 4.2
A&E (2023) 35.82 0.799 0.324 0.250 0.275 0.705 0.782 21.3
D&B (2023) 39.08 0.802 0.323 0.248 0.276 0.713 0.785 18.2
FRAP (ours) 34.22 0.811* 0.327* 0.251* 0.279* 0.750* 0.849* 11.2
5.2 Efficiency Comparisons
Latency values are reported in the last column of Table 1 and also Table 5 in App. B. FRAP enjoys a lower average latency compared to the recent methods, as it does not rely on costly refinement loops. For instance, FRAP is on average $4$ seconds faster than D&B and $5$ seconds faster than A&E for generating one image on the C-CS dataset. Our reported latency measures the average wall-clock time for generating one image on each dataset in seconds with a V100 GPU. Moreover, our method is also more efficient in terms of the number of UNet calls. FRAP consistently use $66$ UNet calls which include $15$ calls for initial selection steps and $51$ calls for adaptive prompt weighting steps. In contrast, A&E and D&B at least perform $76$ UNet calls which include $25$ time-steps with repeated latent code updates each requiring two UNet calls and another $26$ normal time-steps each requiring one UNet call. This number will increase for A&E and D&B with more repeated refinement steps needed on more complex prompts at each time-step.
5.3 Integration with Prompt Rewrite
In addition to the training-free methods, we also consider the recent training-based prompt optimization method Promptist (Hao et al., 2023) which collects an initial dataset and finetunes an LLM to rewrite user prompts. Note that it requires expensive training, and it needs to run the LLM alongside the T2I diffusion model. In Table 3, our results show that although Promptist improves the image aesthetics, it significantly degrades the prompt-image alignment, demonstrated by the lower alignment scores than SD using the original user prompt. FRAP’s adaptive prompt weighting can easily integrate with prompt rewrite methods and could be applied to the rewritten prompt to recover their degraded prompt-image alignment. By applying FRAP on the rewritten prompt of Promptist, we observed improvements in both the prompt-image alignment and overall image quality over the Promptist method as shown in Table 3. In Fig. 2, we provide visual comparisons of FRAP, the LLM prompt rewrite method Promptist (Hao et al., 2023), and the result of combining FRAP and Promptist.
Table 3: Quantitative comparison of SD, training-based LLM prompt optimization/rewrite method Promptist (Hao et al., 2023), applying FRAP on Promptist rewritten prompt, and FRAP, on the complex Color-Object-Scene dataset.
Method
Prompt-Image
Alignment
Overall
Image Quality
Image
Authenticity
Latency (s) $\downarrow$
TTS $\uparrow$
Full-
CLIP $\uparrow$
MOS $\uparrow$
HPS
v2 $\uparrow$
CLIP-
IQA $\uparrow$
CLIP-IQA
Real $\uparrow$
SD (2022) 0.707 0.366 0.254 0.282 0.675 0.770 8.0
Promptist (2023) 0.666 0.358 0.249 0.279 0.698 0.939 9.2
Promptist (2023) + FRAP
0.694 0.377 0.265 0.283 0.702 0.940 15.5
FRAP 0.727 0.381 0.269 0.287 0.676 0.835 14.3
“a pink clock
and a green apple
in the bedroom
“a green backpack and
a yellow chair on the street,
nighttime driving scene
“a red cat and a
black backpack on the street,
nighttime driving scene
“a black cat and
a red suitcase
in the bedroom
Figure 2: Qualitative comparison of SD, training-based LLM prompt rewrite method Promptist (Hao et al., 2023), applying FRAP on Promptist rewritten prompt, and FRAP, on the complex Color-Object-Scene dataset. For each prompt, we show images generated by all four methods (using the same set of seeds). The object tokens are highlighted in green and modifier tokens are highlighted in blue.
5.4 Visual Comparisions
5.4.1 Realistic Appearance and Aesthetics
In Fig. 3, we observe that SD (Rombach et al., 2022) ignores one of the objects (e.g., “crown”, “turtle”, “dog”) in all prompts and often incorrectly binds the color modifiers, yet it maintains a good level of authenticity. Both A&E and D&B generate less realistic images (e.g., “turtle”, “horse”, and “clock” images exhibit cartoonish appearance), likely due to the aforementioned OOD issue. In contrast, FRAP generates images that are more faithful to the prompt, containing all the mentioned objects and well-separated with accurate color bindings (e.g., “clock” and “apple” with correct color bindings instead of a hybrid of two objects as seen for other methods). More importantly, images generated with FRAP maintain a more realistic appearance.